# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ VI-A Autonomous Driving
Real-time processing of complex sensor data is crucial for safe and efficient autonomous driving. Edge-cloud collaboration offers a promising solution by distributing computational workloads between resource-constrained in-vehicle edge devices and powerful cloud servers. Specifically, edge devices can handle time-sensitive tasks like object detection and perception, potentially leveraging lightweight frameworks like Edge YOLO [204], which combines pruned and compressed networks for efficient multi-scale prediction on resource-limited platforms. In addition, the collaborative nature of these systems enables integration of crowdsourced data, enhancing perception capabilities by providing a more comprehensive environmental view, as demonstrated by LiveMap [205, 206], which efficiently processes and integrates multi-vehicle data for real-time dynamic map generation. Beyond perception, LLMs are emerging as powerful tools for motion planning and decision-making [207]. For example, systems like EC-Drive utilize LLMs across both cloud and edge, adapting to data drifts and optimizing communication for reduced latency. Similarly, collaborative policy learning through DRL [205] offers a mechanism for dynamic scheduling and resource allocation, optimizing vehicle scheduling and offloading decisions for improved system robustness and adaptability. Adaptive edge-cloud collaboration, aided by large vision models like in LAECIPS [208], further enhances performance and adaptability of IoT-based perception systems through efficient hard input mining. Concurrently, the fusion of DRL and edge computing facilitates real-time monitoring and control optimization in IoT environments [21], enabling predictive system state analysis and dynamic resource allocation for enhanced control.
@@ VI-B Smart Healthcare
AI-driven collaborative computing transforms smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig. 6. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [209]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [210]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [211], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [28]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [212]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [213]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [214]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [215].
@@ VI-C Industrial Automation
Intelligent automation systems benefit significantly from collaborative computing architectures that deliver real-time fault diagnosis and predictive maintenance capabilities [21]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [80], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [216]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [65].
@@ VI-D Smart Cities
Urban infrastructure optimization leverages collaborative computing to enhance traffic management and environmental monitoring capabilities [217, 142]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [218]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [16, 219, 220]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.
@@ VII-A1 System Performance
Throughput, latency, and scalability are crucial metrics for evaluating the effectiveness and responsiveness of ECCC systems [116, 20, 221]. Throughput measures the volume of data processed or tasks completed within a given timeframe, with higher throughput indicating efficient handling of substantial workloads [222]. Conversely, latency quantifies the delay between request initiation and completion; minimizing latency is paramount for real-time and interactive applications where responsiveness is critical [221]. Scalability, in turn, refers to a system’s ability to adapt to increasing workloads or expanding infrastructure without performance degradation, accommodating growth in user demands or data volume while maintaining acceptable throughput and latency [203]. The distributed and heterogeneous nature of edge-cloud environments amplifies the significance of these metrics, as the interplay between edge and cloud resources necessitates careful optimization. For example, the study in [116] benchmarks a container-based edge computing system, demonstrating the impact of data synchronization on scalability across different deployment strategies. Similarly, the SPEC-RG reference architecture in [20] provides a framework for analyzing resource management and performance, while [221] highlights the potential for performance inversion due to edge queuing delays. Evaluating these metrics is essential for understanding system efficiency and informing design choices for optimized performance in resource-constrained and dynamic edge-cloud environments [62].
@@ VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [223], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [224, 225]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [226], necessitating compression techniques like pruning, quantization, or knowledge distillation [70]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [226]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [20]. Efficiency optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [227].
@@ VII-A3 Quality of Service
Evaluating the performance of edge-cloud collaborative systems necessitates an analysis of QoS metrics, which directly reflect the system’s capacity to fulfill user expectations and provide consistent, dependable service [228]. Key QoS metrics include availability, reliability, and user satisfaction. Availability quantifies the proportion of time a system remains operational and accessible, a crucial aspect in edge-cloud systems where interruptions significantly impact user experience and application performance [20]. High availability relies on robust system design, redundant infrastructure, and effective fault tolerance mechanisms, further enhanced by dynamic resource management and adaptive optimization techniques [229]. Reliability, in turn, assesses the system’s ability to consistently deliver the expected service level over time, encompassing fault tolerance, error recovery, and performance maintenance under varying workloads [230]. Dependency-aware offloading and collaborative resource management contribute to improved reliability by optimizing task placement and resource allocation. While often subjective, user satisfaction remains a crucial QoS indicator, reflecting the overall user experience, including performance, ease of use, and perceived value [231]. In edge-cloud systems, factors such as latency, throughput, and the quality of AI-driven services influence user satisfaction. Specifically, QoE serves as a proxy for user satisfaction, directly linked to QoS parameters like latency and loss in applications like video streaming [232]. Similarly, in industrial automation, the relationship between QoS and QoE is critical for optimizing network resource management and ensuring customer satisfaction.
@@ VII-A4 Security and Privacy
In edge-cloud collaborative systems, robust security and privacy are crucial for maintaining trust and protecting sensitive information, especially given the sensitive nature of processed data [233]. Therefore, evaluating these systems requires metrics focusing on data breaches, encryption strength, and regulatory compliance. Data breach frequency directly indicates system vulnerability by quantifying unauthorized accesses within a given timeframe [198]. A lower frequency suggests stronger security, and analyzing breach characteristics can inform improved protocols, potentially leveraging frameworks like the transparency, accountability, and blockchain framework [233]. Encryption strength assesses cryptographic robustness, including key length, algorithms, and resistance to attacks. Strong encryption is paramount in distributed edge-cloud environments [194], and combining techniques like lightweight symmetric encryption and ciphertext-policy attribute-based encryption can enhance security [194]. Additionally, compliance with regulations like GDPR and CCPA is essential. Metrics in this area assess adherence to data anonymization, user consent, and data retention policies [234]. Evaluating privacy protectability, along with techniques like PrivaScissors, can further enhance privacy protection [235, 182].
@@ VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems requires standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs and optimization strategies.
TABLE VI: Comparison of edge-cloud benchmarking tools and platforms.
Tool Focus Area Key Features Advantages Limitations Use Case Scenarios
Edge AIBench [236] End-to-end edge AI performance Models typical edge scenarios (ICU monitoring, autonomous vehicles), comprehensive client-edge-cloud evaluation Realistic application scenarios Comprehensive performance metrics Multi-tier evaluation coverage Limited customization options Fixed scenario constraints Domain-specific focus Healthcare monitoring, autonomous driving, real-time inference applications
EdgeBench [237] Workflow-based edge computing Customizable workflow logic, configurable data storage, flexible stage distribution across computing tiers High customization flexibility Workflow-centric approach Adaptable to diverse applications Complex configuration process Requires workflow expertise Limited standardized scenarios Custom application workflows, multi-stage processing pipelines, research environments
BenchFaaS [238] Serverless edge computing Heterogeneous testbeds (VMs, Raspberry Pi, cloud platforms), FaaS performance evaluation Serverless focus Heterogeneous platform support Function-level granularity Limited to FaaS paradigm Cold start overhead Platform-dependent results Serverless applications, function-as-a-service deployment, microservices evaluation
KheOps [239] Reproducible edge-to-cloud experiments Collaborative environment, experiment repository, notebook interface, multi-platform methodology Strong reproducibility support Collaborative features Multi-platform compatibility Learning curve for new users Resource-intensive setup Limited real-time capabilities Research collaboration, experimental reproducibility, academic studies
OpenPerf [240] Open-source project benchmarking General performance evaluation framework, standardized metrics, community-driven development Open-source transparency Community contributions Standardized approach Not edge-cloud specific Generic rather than specialized Limited edge device support General performance testing, open-source projects, baseline comparisons
Table VI provides a comprehensive comparison of existing benchmarking tools and their respective strengths and limitations across different edge-cloud deployment scenarios. Standardized datasets and workloads, discussed in Sec. VII-B1, provide a common basis for evaluation, minimizing variability stemming from data discrepancies. Well-defined testing environments, explored in Sec. VII-B2, further enhance reliability and reproducibility by controlling hardware and software configurations, network setups, and other influential environmental factors. Performance comparison frameworks and evaluation methodologies, including simulation, emulation, and real-world testing approaches, offer structured methodologies and tools for identifying performance bottlenecks and areas for improvement in edge-cloud systems, with detailed analysis provided in the Supplementary Material.
The choice of appropriate benchmarking tools depends on several factors including application domain, system architecture, and evaluation objectives. Edge AIBench [236] excels in scenarios requiring realistic application modeling with comprehensive end-to-end performance analysis, making it particularly suitable for healthcare and autonomous driving applications where domain-specific workloads are crucial. EdgeBench [237] offers superior flexibility for research environments and custom applications that require tailored workflow configurations, though it demands significant expertise in workflow design and configuration. BenchFaaS [238] provides specialized evaluation capabilities for serverless edge computing paradigms, offering detailed insights into function-level performance characteristics but with limited applicability outside the FaaS domain. KheOps [239] stands out for collaborative research environments where reproducibility and experiment sharing are paramount, though it may require substantial initial setup overhead. The selection process should consider factors such as workload representativeness, customization requirements, evaluation scope, and available expertise. Additionally, the benchmarking tool should align with the specific performance metrics of interest, whether focusing on latency, throughput, energy efficiency, or comprehensive system-level analysis.
@@ VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [241]. Unified standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [236] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [237] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [20]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [239] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. The availability of standard datasets and workloads also facilitates the development of new optimization techniques [242], allowing researchers to evaluate their methods on established benchmarks [227].

### v5

@@ VI-A Autonomous Driving
Real-time processing of complex sensor data is crucial for safe and efficient autonomous driving. Edge-cloud collaboration offers a promising solution by distributing computational workloads between resource-constrained in-vehicle edge devices and powerful cloud servers. Specifically, edge devices can handle time-sensitive tasks like object detection and perception, potentially leveraging lightweight frameworks like Edge YOLO [liang2022edge], which combines pruned and compressed networks for efficient multi-scale prediction on resource-limited platforms. In addition, the collaborative nature of these systems enables integration of crowdsourced data, enhancing perception capabilities by providing a more comprehensive environmental view, as demonstrated by LiveMap [liu2021livemap, liu2023realtime], which efficiently processes and integrates multi-vehicle data for real-time dynamic map generation. Beyond perception, LLMs are emerging as powerful tools for motion planning and decision-making [chen2024edgecloud]. For example, systems like EC-Drive utilize LLMs across both cloud and edge, adapting to data drifts and optimizing communication for reduced latency. Similarly, collaborative policy learning through DRL [liu2021livemap] offers a mechanism for dynamic scheduling and resource allocation, optimizing vehicle scheduling and offloading decisions for improved system robustness and adaptability. Adaptive edge-cloud collaboration, aided by large vision models like in LAECIPS [hu2024laecips], further enhances performance and adaptability of IoT-based perception systems through efficient hard input mining. Concurrently, the fusion of DRL and edge computing facilitates real-time monitoring and control optimization in IoT environments [xu2024fusion], enabling predictive system state analysis and dynamic resource allocation for enhanced control.
@@ VI-B Smart Healthcare
AI-driven collaborative computing transforms smart healthcare systems into efficient, secure, and personalized ecosystems, as shown in Fig. 9. We explore how AI-driven edge-cloud collaboration revolutionizes healthcare delivery and improves patient outcomes by investigating FL for medical image analysis, edge-cloud platforms for remote patient monitoring, and resource allocation optimization for Internet of medical things (IoMT) devices. FL offers a promising technique for training AI models on distributed medical datasets without compromising patient privacy, allowing collaborative model training across multiple healthcare providers while keeping sensitive data localized [shaik2023remote]. Specifically, FL enables robust diagnostic model development, including COVID-19 diagnosis using multimodal data like X-rays and ultrasound images [qayyum2022collaborative]. Edge-cloud platforms are crucial for enabling remote patient monitoring and personalized healthcare [wu2020personalized], leveraging edge device computation and cloud storage for real-time monitoring and personalized interventions. For example, wearable sensors and IoMT devices collect patient data, process it at the edge for immediate feedback, and transmit it to the cloud for analysis and storage [kanduri2024edgecentric]. AI algorithms deployed on edge devices analyze this data, providing personalized recommendations and improving intervention effectiveness [shaik2022fedstack]. Efficient resource allocation is also essential for optimizing IoMT device performance and ensuring timely healthcare delivery [nguyen2022federated]. Consequently, AI-driven task offloading and resource management dynamically allocate tasks and resources between edge and cloud based on real-time demands and availability, minimizing latency and improving system efficiency. Hierarchical FL approaches further enable multi-party collaboration and efficient model training for anomaly detection in IoMT data, leveraging digital twins and edge cloudlets for real-time patient health insights and timely interventions [gupta2021hierarchical]. Integrating FL within the healthcare Metaverse enhances these capabilities by improving privacy, scalability, and interoperability [bashir2023survey].
@@ VI-C Industrial Automation
Intelligent automation systems benefit significantly from collaborative computing architectures that deliver real-time fault diagnosis and predictive maintenance capabilities [xu2024fusion]. Leveraging both cloud computational power and edge real-time processing capabilities, industrial automation systems achieve greater agility through optimized production processes and resource allocation [zhang2024cloud], where cloud-based platforms analyze historical data to generate optimized production schedules while edge devices adapt these schedules in real-time based on local conditions. Distributed intelligence across multiple edge devices handles complex manufacturing tasks more efficiently and enhances fault tolerance since individual device failures do not necessarily compromise the entire system [zeng2021coedge]. As edge intelligence, anomaly detection, cloud-edge collaboration, and distributed intelligence converge, industrial automation systems evolve toward more sophisticated and autonomous manufacturing processes [tang2024anomaly].
@@ VI-D Smart Cities
Urban infrastructure optimization leverages collaborative computing to enhance traffic management and environmental monitoring capabilities [firdaus2023joint, arroba2024sustainable]. Edge devices collect and analyze real-time data for adaptive traffic control and congestion reduction, while distributed sensor networks monitor environmental factors like air quality and noise levels. Subsequently, cloud platforms aggregate and analyze this data to inform urban planning and environmental protection. In addition, AI-driven resource allocation optimizes smart grid energy consumption [taik2020electrical]. Specifically, edge devices monitor energy usage, and cloud-based algorithms dynamically adjust energy distribution to minimize waste and improve grid stability. FL is crucial for privacy-preserving data analysis and decision-making in these applications [zhou2024enhancing, zhu2024flight, feng2024leveraging]. Through FL, stakeholders such as government agencies and private companies can collaboratively train models on combined data without sharing sensitive information, leading to more accurate and robust models for urban planning and public safety.
@@ VII-A1 System Performance
Throughput, latency, and scalability are crucial metrics for evaluating the effectiveness and responsiveness of ECCC systems [carpio2020engineering, jansen2023specrg, ali-eldin2021hidden]. Throughput measures the volume of data processed or tasks completed within a given timeframe, with higher throughput indicating efficient handling of substantial workloads [michalke2023scaling]. Conversely, latency quantifies the delay between request initiation and completion; minimizing latency is paramount for real-time and interactive applications where responsiveness is critical [ali-eldin2021hidden]. Scalability, in turn, refers to a system’s ability to adapt to increasing workloads or expanding infrastructure without performance degradation, accommodating growth in user demands or data volume while maintaining acceptable throughput and latency [queralta2021blockchain]. Given that the distributed and heterogeneous nature of edge-cloud environments amplifies the significance of these metrics, the interplay between edge and cloud resources necessitates careful optimization. For example, the study in [carpio2020engineering] benchmarks a container-based edge computing system, demonstrating the impact of data synchronization on scalability across different deployment strategies. Similarly, SPEC-RG [jansen2023specrg] provides a reference framework for analyzing resource management and performance, while [ali-eldin2021hidden] highlights the potential for performance inversion due to edge queuing delays. Evaluating these metrics is essential for understanding system efficiency and informing design choices for optimized performance in resource-constrained and dynamic edge-cloud environments [rublein2024improved].
@@ VII-A2 Model Efficiency
Efficient AI models are essential for effective deployment and operation within the resource-constrained edge-cloud continuum. Key metrics for quantifying model efficiency include computational cost, memory usage, and inference time, informing model selection and deployment strategies. Computational cost, typically measured in FLOPS [caspart2022precise], quantifies processing demands. Higher FLOPS values indicate greater computational requirements and potentially increased energy consumption, impacting resource-limited edge devices and overall system sustainability [dehghani2022efficiency, richins2021ai]. Memory usage reflects a model’s storage footprint; large models may exceed the limited capacity of edge devices [singh2025impact], necessitating compression techniques like pruning, quantization, or knowledge distillation [wu2022serving]. Inference time, the duration for processing input and generating output, directly impacts application responsiveness [singh2025impact]. In latency-sensitive applications like autonomous driving or real-time video analytics, minimizing inference time is crucial [jansen2023specrg]. Efficiency optimization often involves balancing model accuracy and computational efficiency, potentially employing strategies like model partitioning across different devices [rosendo2022distributed].
@@ VII-A3 Quality of Service
Evaluating the performance of edge-cloud collaborative systems necessitates an analysis of QoS metrics, which directly reflect the system’s capacity to fulfill user expectations and provide consistent, dependable service [zeng2024framework]. Key QoS metrics include availability, reliability, and user satisfaction. Availability quantifies the proportion of time a system remains operational and accessible, a crucial aspect in edge-cloud systems where interruptions significantly impact user experience and application performance [jansen2023specrg]. High availability relies on robust system design, redundant infrastructure, and effective fault tolerance mechanisms, further enhanced by dynamic resource management and adaptive optimization techniques [sedlak2024equilibrium]. Reliability, in turn, assesses the system’s ability to consistently deliver the expected service level over time, encompassing fault tolerance, error recovery, and performance maintenance under varying workloads [simon2021parsimonious]. Dependency-aware offloading and collaborative resource management contribute to improved reliability by optimizing task placement and resource allocation. While often subjective, user satisfaction remains a crucial QoS indicator, reflecting the overall user experience, including performance, ease of use, and perceived value [arnab2021analysis]. In edge-cloud systems, factors such as latency, throughput, and the quality of AI-driven services influence user satisfaction. Specifically, QoE serves as a proxy for user satisfaction, directly linked to QoS parameters like latency and loss in applications like video streaming [vidacs2023network]. Similarly, in industrial automation, the relationship between QoS and QoE is critical for optimizing network resource management and ensuring customer satisfaction.
@@ VII-A4 Security and Privacy
In edge-cloud collaborative systems, robust security and privacy are crucial for maintaining trust and protecting sensitive information, especially given the sensitive nature of processed data [xu2023blockchainbased]. Therefore, evaluating these systems requires metrics focusing on data breaches, encryption strength, and regulatory compliance. Data breach frequency directly indicates system vulnerability by quantifying unauthorized accesses within a given timeframe [zobaed2022aidriven]. A lower frequency suggests stronger security, and analyzing breach characteristics can inform improved protocols, potentially leveraging frameworks like the transparency, accountability, and blockchain framework [xu2023blockchainbased]. Encryption strength assesses cryptographic robustness, including key length, algorithms, and resistance to attacks. Strong encryption is paramount in distributed edge-cloud environments [yao2024efficient], and combining techniques like lightweight symmetric encryption and ciphertext-policy attribute-based encryption can enhance security [yao2024efficient]. Additionally, compliance with regulations like GDPR and CCPA is essential. Metrics in this area assess adherence to data anonymization, user consent, and data retention policies [zhao2022using]. Evaluating privacy protectability, along with techniques like PrivaScissors, can further enhance privacy protection [shi2023privacy, duan2023privascissors].
@@ VII-B Benchmarking Methods
Performance benchmarking of edge-cloud collaborative systems necessitates standardized evaluation approaches to facilitate meaningful comparisons across diverse system designs, with a detailed analysis of tools provided in the Supplementary Material (see LABEL:t7-benchmark-comparison in LABEL:sec:supp-benchmarking). Essential to this process are standardized datasets and controlled testing environments that minimize variability, alongside comprehensive comparison frameworks employing simulation, emulation, and real-world testing to identify bottlenecks. For instance, selecting appropriate tools depends on application domains and evaluation objectives, Edge AIBench [hao2019edge] provides realistic application modeling for healthcare and autonomous driving, while EdgeBench [yang2024edgebench] offers flexibility for custom workflow configurations. Moreover, BenchFaaS [carpio2023benchfaas] specializes in serverless paradigms, whereas KheOps [rosendo2023kheops] facilitates collaborative research reproducibility, ultimately requiring that the optimal choice aligns workload representativeness and performance metrics with specific system requirements.
@@ VII-B1 Standard Datasets and Workloads
Consistent and comparable performance evaluation across different studies of ECCC systems requires the use of standard datasets and workloads [mohammadi2018continuous]. Unified standardization clarifies the strengths and weaknesses of various architectures, optimization techniques, and algorithms, enabling researchers to build upon existing work and accelerate the development of innovative solutions. For example, Edge AIBench [hao2019edge] offers benchmarks modeling typical edge scenarios like ICU patient monitoring and autonomous vehicles, focusing on data distribution and workload collaboration across client devices, the edge, and the cloud for comprehensive end-to-end performance evaluation. Similarly, EdgeBench [yang2024edgebench] provides a workflow-based benchmarking approach, allowing customization of workflow logic, data storage, and the distribution of workflow stages across computing tiers to tailor benchmarks to specific application scenarios. Appropriate workload selection is also crucial [jansen2023specrg]. Specifically, workloads should reflect real-world application characteristics, including data intensity, computational complexity, and communication patterns, particularly in distributed intelligence scenarios where data and model distribution, as well as communication overhead for model training and inference, are critical factors. KheOps [rosendo2023kheops] further emphasizes reproducibility in edge-to-cloud experiments by providing a collaborative environment with an experiment repository, a notebook environment, and a multi-platform methodology, promoting transparency and rigor in performance evaluation. Moreover, the availability of standard datasets and workloads facilitates the development of new optimization techniques [zhang2022ents], allowing researchers to evaluate their methods on established benchmarks [rosendo2022distributed].